OpenStudy (anonymous):

if a2=-4 and a5=-3, find the first and sixth term of this arithmetic sequence.

OpenStudy (jennychan12):

the formula for an arith. sequence is an = a1+(n-1)d so... adjust it accordingly, so that u get... a5 = a2 + 3d Find d. Then find a1 using another adjusted formula that is, a5 = a1 + 4d (plug in d from above formula) hope it helps?